### v2

# CamReasoner: Reinforcing Camera Movement Understanding via Structured Spatial Reasoning

Abstract
/Understanding camera dynamics is a fundamental pillar of video spatial intelligence. However, existing multimodal models predominantly treat this task as a black-box classification, often confusing physically distinct motions by relying on superficial visual patterns rather than geometric cues. We present CamReasoner, a framework that reformulates camera movement understanding as a structured inference process to bridge the gap between perception and cinematic logic. Our approach centers on the Observation-Thinking-Answer (O-T-A) paradigm, which compels the model to decode spatio-temporal cues such as trajectories and view frustums within an explicit reasoning block. To instill this capability, we construct a Large-scale Inference Trajectory Suite comprising 18k SFT reasoning chains and 38k RL feedback samples. Notably, we are the first to employ RL for logical alignment in this domain, ensuring motion inferences are grounded in physical geometry rather than contextual guesswork. By applying Reinforcement Learning to the Observation-Think-Answer (O-T-A) reasoning paradigm, CamReasoner effectively suppresses hallucinations and achieves state-of-the-art performance across multiple benchmarks.
[…]
1 Introduction
Camera movement serves as the invisible brushstroke of cinematic storytelling, acting as a vital conduit to convey subtle emotions and orchestrate audience immersion. Whether through Scorsese’s seamless tracking shots that create a seductive flow, or Spielberg’s dynamic dollies that heighten suspense and awe, these trajectories represent the intentionality of a director shaping the viewer’s perception of the physical and emotional world. In the landscape of Multimodal Large Language Models (MLLMs), camera movement understanding has emerged as a critical frontier that bridges the gap between passive pixel perception and active semantic reasoning. From the perspective of understanding, mastering these dynamics allows MLLMs to effectively decouple complex object movements from the observer’s ego-motion, even when camera intrinsic parameters are unknown, thereby enabling a more robust comprehension of 3D spatial geometry and narrative intent. This transition from surface-level recognition to deep structural analysis transforms an AI from a mere observer into a cinematic reasoner capable of interpreting the "why" behind every frame. Furthermore, from the perspective of generation, camera movement acts as the essential cornerstone for controllable video synthesis. As the frontier of AI shifts toward purposeful visual artistry, the ability to generate physically consistent videos relies fundamentally on the model’s grasp of how trajectories interact with scene depth and perspective. By integrating these motion priors, MLLMs can move beyond disorganized pixel synthesis toward achieving true world modeling capabilities, where every shift in perspective is both physically grounded and narratively coherent.
Despite the growing demand for such sophisticated motion perception, current methodologies for camera movement understanding remain split between two constrained paradigms that struggle to meet these requirements. Traditional geometric approaches, such as SfM [schonberger2016structure, wang2024vggsfm, zhang2023protein] and SLAM [davison2007monoslam, Engel2014LSD], rely on per-frame camera pose estimation to reconstruct trajectories, yet often require prior knowledge of camera intrinsics to maintain accuracy. These systems are fragile in dynamic environments because they often fail to disentangle the camera’s ego-motion from the complex movements of subjects within the scene. In contrast, modern learning-based Video-Language Models (VLMs) [comanici2025gemini, qwen3vl, hurst2024gpt] offer a promising alternative by treating motion understanding as a semantic task. Leveraging vast pre-training corpora, these models achieve impressive alignment between visual cues and cinematic terminology, while even demonstrating an emerging capacity for high-level reasoning [liu2025shotbench]. However, these models predominantly operate as black boxes that leap from pixels to labels without explicit reasoning. This shortcut often triggers hallucinations because the model relies on contextual cues like a running athlete instead of decoding the actual physics of the lens. Consequently, they struggle to distinguish physically distinct yet visually similar operations such as a zoom versus a dolly. These limitations necessitate a framework that bridges the gap between raw perception and structured cinematic logic.
Figure 1: Overview of CamReasoner. We propose the <observation>-<think>-<answer>paradigm to improve camera movement understanding. The figure illustrates our model generating detailed visual observations and logical thinking for movements like truck left and pan right.
To bridge this gap, we introduce CamReasoner, a framework that reformulates camera movement understanding as a structured ego-motion inference task. Central to our approach is the Observation-Thinking-Answer (O-T-A) paradigm, which we implement through a sequential training pipeline of Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL). To support this process, we constructed a Large-scale Inference Trajectory Suite comprising 18k SFT and 38k RL samples. While the SFT data transforms static labels into explainable reasoning chains to instill cinematic logic, the RL samples enforce logical alignment to ground final verdicts in 3D spatio-temporal physics. By explicitly tracing trajectories and analyzing view frustums, our model establishes a rigorous logical premise before delivering a verdict, effectively eliminating contextual guesswork. As the first to apply Reinforcement Learning for logical alignment in this domain, CamReasoner-7B achieves state-of-the-art results, reaching 78.4% in binary classification and 74.5% in VQA. Notably, it excels in challenging Confusable Motion scenarios (60.7%) by distinguishing physical movements like truck versus pan through geometric reasoning. By integrating structural motion priors, CamReasoner provides a scalable path for MLLMs to achieve precise cinematic control and physically-consistent world modeling.
Our main contributions are summarized as follows:
  • Structured Reasoning Paradigm: We propose the Observation-Thinking-Answer (O-T-A) paradigm, reformulating camera movement understanding from black-box classification into a structured ego-motion inference task that decouples camera trajectories from dynamic scenes.
  • Inference Trajectory Suite: We construct a Large-scale Inference Trajectory Suite with 18k SFT and 38k RL samples, transforming static labels into dense reasoning chains to instill spatio-temporal cinematic logic into MLLMs.
  • RL-driven Logical Alignment: We are the first to employ Reinforcement Learning for logical alignment in this domain, achieving state-of-the-art 78.4% and 74.5% accuracy in binary classification and visual question answering tasks.
2 Related Work
2.1 Camera Movement Understanding
Camera movement understanding bridges geometric perception and semantic reasoning, yet current methods face significant bottlenecks. Traditional geometric approaches rely on SfM-based pseudo ground-truth, which is largely restricted to static scenes and ignores scene context [zhou2018stereo, ling2024dl3dv, jin2024stereo4d, schonberger2016structure]. Categorical recognition tasks, while providing semantic labels, often suffer from poor specification—confusing rotation with translation (e.g., pan vs. truck) and failing to capture co-occurring motions in complex cinematography [huang2020movienet, rao2020shot, argaw2022anatomy]. Although recent benchmarks like ShotBench [liu2025shotbench, wu2025refineshot] evaluate nuanced cinematic grammar, existing methodologies lack a structured reasoning process to explain the visual cues driving their interpretations [hong2025motion, wang2024tarsier]. Consequently, there remains a critical need for a framework that integrates spatial evidence with logical deduction to achieve a holistic understanding across diverse video domains [zhou2018stereo, huang2020movienet, argaw2022anatomy, liu2025shotbench].
2.2 Multimodal Reinforcement Learning
Inspired by the reasoning success of LLMs, recent works have extended structured thinking to MLLMs across diverse visual tasks [li2025star, sun2025reinforcement, feng2025video, sun2025spacevista, zhou2025reinforced, duan2025codeplot, chen2025advancing, chen2025ares, meng2025open]. These include task-specific reasoning models for image VQA [huang2025vision, mei2025surveycontextengineeringlarge, mei2025a1], video understanding [feng2025video, li2025videochat], object detection [yu2025perception, shen2025vlm], segmentation [you2025seg], and temporal grounding [wang2025time]. Notably, frameworks like SophiaVL-R1 [fan2025sophiavl], OneThinker [feng2025onethinker], and FrameMind [ge2025framemind] introduce unified reasoning via reinforcement learning, while Thinking-with-Sound [xiong2025thinking] extends this paradigm to the audio domain through acoustic tool manipulation. However, despite the surge in RL-based multimodal reasoning, its application to the specialized domain of camera movement understanding remains unexplored.
3 Method
We introduce a two-stage training strategy combining an SFT cold start with Group Relative Policy Optimization (GRPO) [shao2024deepseek] to instill the O-T-A reasoning paradigm into camera movement understanding. This pipeline ensures the model efficiently acquires structured reasoning traces while maintaining format adherence and task accuracy. To support this, we curate distinct datasets specifically tailored for sequential Supervised Fine-Tuning and Reinforcement Learning phases.
3.1 Data Collection and Curation
High-quality, structured data is pivotal for teaching the model cinematic logic. We constructed two specialized datasets to facilitate the O-T-A paradigm across the SFT and RL stages.
3.1.1 CamReasoning-SFT-18k Construction.
To instill structured reasoning, we curate the CamReasoning-SFT-18k dataset via an answer-conditioned synthesis pipeline. As shown in Fig. 2, we utilize Qwen2.5-VL-72B to generate 38,672 initial trajectories from original video-QA pairs, enforcing the <observation>, <think>, and <answer> format. To ensure logical coherence, we implement a multi-dimensional filtration process using Qwen3 [qwen3vl] to verify samples based on format adherence, label accuracy, and reasoning consistency. This results in 18,541 high-fidelity samples that serve as cold-start data for teaching the model step-by-step cinematic inference.
3.1.2 CamReasoning-RL-38k
For the reinforcement learning stage, we construct the CamReasoning-RL-38k dataset to drive GRPO [shao2024deepseek]. Derived from the CameraBench training split, this dataset provides a diverse array of camera movement scenarios converted into our standardized instruction schema. Unlike the SFT phase, these samples lack reasoning annotations; instead, they serve as environmental queries for the model to refine its self-generated reasoning paths. This setup allows the model to optimize logical robustness and prediction accuracy through trial and error, guided by task-specific reward signals.
Figure 2: Data generation pipeline for CamReasoning-SFT-18k. We utilize an answer-conditioned process to generate 38,672 initial trajectories, which are then filtered for format, accuracy, and consistency to retain 18,541 high-quality samples.
3.2 Training Strategy
Our training pipeline comprises two progressive stages designed to endow the model with robust O-T-A reasoning capabilities for camera movement understanding. We first employ SFT as a cold-start initialization to adapt the model to the structured <observation>-<think>-<answer>format. Subsequently, we advance to a RL phase utilizing GRPO to further refine the reasoning logic and align outputs with task requirements. Crucially, to mitigate the optimization instability inherent in single-task RL, we incorporate the EMA-GRPO strategy for robust advantage normalization.
3.2.1 Supervised Fine-tuning
We initiate the training process with SFT to bootstrap the model’s capability in generating structured O-T-A (Observation-Think-Answer) reasoning traces. As pre-trained MLLMs lack exposure to our specialized reasoning format encompassing <observation>, <think>, and <answer> tags, SFT provides the essential cold-start initialization. This phase instills adherence to the required formatting and logic patterns, ensuring the model can generate valid reasoning chains and camera movement predictions as defined in our dataset.
Formally, the SFT objective is to minimize the negative log-likelihood of the target sequence $o$ (which explicitly concatenates the observation, reasoning trace, and final answer). Given the input video $v$ and query $q$, the loss function is formulated as:
$\mathcal{L}_{\text{SFT}}(\theta)=-\mathbb{E}_{(v,q,o)\sim\mathcal{D}_{\text{SFT}}}\left[\sum_{t=1}^{|o|}\log\pi_{\theta}\left(o_{t}\mid v,q,o_{<t}\right)\right],$ (1)
where $o_{t}$ represents the $t$-th token of the target sequence $o$, and $o_{<t}$ denotes the preceding tokens. $\pi_{\theta}$ is the policy of the multimodal model parameterized by $\theta$. This stage enables the model to predict each subsequent token conditioned on the visual input, the query, and the generation history, establishing a robust foundation for the subsequent reinforcement learning stage.
3.2.2 Reinforcement Learning Framework
To further align the model’s reasoning capabilities with task requirements and ensure robust logical chains, we employ RL. The training framework is structured into the following components:
Problem Formulation.
We formulate the camera movement understanding task as a sequential decision-making process. Given an input video clip $v$ and a textual query $q$ regarding the camera motion, the multimodal policy $\pi_{\theta}$ generates a structured response sequence $o$. This output $o$ comprises the visual observation, the explicit reasoning trace, and the final predicted camera movement category. Our optimization objective is to maximize the expected reward of these generated sequences under the policy:
$J(\theta)=\mathbb{E}_{o\sim\pi_{\theta}(\cdot|v,q)}[R(o)].$ (2)
Reward Function Design.
Designing an effective reward function is critical for guiding the O-T-A reasoning process. We define a composite reward $R(o)$ that evaluates both the structural validity and the factual correctness of the output. The reward consists of two components: (1) Format Reward $r_{fmt}$: To enforce the strict reasoning structure, we apply a regular expression check. The model receives a reward of $1.0$ if and only if the generated output strictly adheres to the sequence of <observation>, <think>, and <answer> tags; otherwise, $r_{fmt}=0.0$. (2) Accuracy Reward $r_{acc}$: Conditional on passing the format check, we extract the prediction enclosed within the <answer> tags. If the predicted content matches the ground truth camera movement label, $r_{acc}$ is set to a positive score (e.g., $1.0$); otherwise, a penalty is applied. If the format check fails, $r_{acc}$ is explicitly set to $0.0$.
The final reward $R(o)$ is a weighted interpolation between accuracy and format adherence. Let $\lambda$ denote the format weight. The overall reward is formulated as:
$R(o)=(1-\lambda)\cdot r_{acc}+\lambda\cdot r_{fmt}.$ (3)
This formulation balances the trade-off between structure adherence and prediction accuracy, ensuring the model learns to generate valid reasoning traces while optimizing for correct answers.
GRPO.
To optimize the policy efficiently without the memory overhead of a value network, we utilize the GRPO algorithm. For each query $q$, we sample a group of $G$ outputs $\{o_{1},o_{2},\dots,o_{G}\}$ from the current policy $\pi_{\theta_{old}}$. GRPO estimates the baseline using the group average, avoiding the need for a separate critic model. The training objective is to minimize the following surrogate loss: Let $r_{i}(\theta)=\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}$ be the probability ratio.
$\mathcal{L}_{GRPO}(\theta)$ $=-\frac{1}{G}\sum_{i=1}^{G}\Bigg[\min\Big(r_{i}(\theta)A_{i},$ (4)
$\quad\text{clip}\big(r_{i}(\theta),1-\epsilon,1+\epsilon\big)A_{i}\Big)-\beta D_{KL}\Bigg],$
where $\epsilon$ is the clipping parameter, $A_{i}$ is the advantage value, and $\beta$ controls the strength of the KL divergence penalty $D_{KL}$ (measured against the reference model $\pi_{ref}$).
Stability Strategy: EMA-GRPO.
To ensure training stability, we adopt the EMA-GRPO [feng2025onethinker] strategy for advantage normalization. Standard GRPO normalizes rewards using the immediate batch standard deviation. This introduces significant gradient fluctuations, as low-variance rollouts can lead to artificially inflated advantages and unstable updates. To mitigate this, we employ an exponential moving average (EMA) to estimate a globally stable standard deviation, denoted as $\sigma(t)$. The advantage is computed as:
$A_{i}(t)=\frac{R_{i}-\mathrm{mean}(\{R_{j}\})}{\sigma(t)}.$ (5)
Here, $\sigma(t)$ captures the intrinsic reward scale and filters out high-frequency noise from individual batches. EMA-GRPO effectively enforces a globally consistent trust region by stabilizing the reward scale, preventing spurious large policy updates caused by accidental low-entropy rollout groups. This approach decouples gradient scaling from immediate rollout randomness and ensures robust convergence during single-task training.
4 Experiments
4.1 Setup
Evaluation Settings
We evaluate our model using data from the CameraBench [lin2025towards] dataset, focusing on binary classification of motion primitives and the 10K-sample VQA subset covering 9 skills and 81 sub-tasks. We tested the accuracy of multiple existing MLLMs using the questions and videos across the datasets. To ensure visual grounding, each question is paired with two videos yielding opposite answers. During inference, we standardize video input to a maximum of 64 frames and 262,144 pixels ($128\times 32\times 32$) sampled at 1 FPS, with a batch size of 128. This setup evaluates the model across object-centric motion, scene dynamics, and complex logical verification.
Implementation Details
The RL phase is implemented using the verl framework with a learning rate of $2\times 10^{-6}$ and bf16 mixed-precision. We employ a rollout batch size of 128 and a global batch size of 32. To ensure stability, we incorporate an online filtering mechanism that prunes samples in the top and bottom 1% of the reward distribution. The total reward is defined as $R=(1-\lambda)\cdot r_{acc}+\lambda\cdot r_{fmt}$, with $\lambda=0.1$. Crucially, the RL policy is initialized from the first SFT epoch to provide a foundational reasoning grasp while maintaining sufficient entropy for exploration.
4.2 Quantitative Results
Table 1: Binary classification performance on camera movement understanding. CamReasoner-7B achieves state-of-the-art results across most categories, significantly outperforming SfM/SLAM baselines and general-purpose MLLMs.
{NiceTabular}
lccccccccccccccc|$c$ Model Translation (Dolly/Pedestal/Truck) Zooming Rotation (Pan/Tilt/Roll) Static Avg
In Out Up Down Right Left In Out Right Left Up Down CW CCW
Random Chance 29.3 9.7 6.7 8.6 15.8 11.5 11.1 10.2 15.0 15.4 12.7 7.7 8.9 10.2 9.7 12.2
SfM/SLAM
COLMAP [schonberger2016structure] 36.2 13.1 11.9 19.7 34.1 30.0 13.9 14.2 43.9 46.4 28.3 19.1 42.1 48.7 7.5 27.3
VGGSFM [wang2024vggsfm] 56.6 28.9 28.7 38.2 48.9 35.3 21.7 17.3 60.9 58.7 46.6 43.3 61.4 55.5 16.7 41.3
DUSt3R [wang2024dust3r] 58.9 24.0 30.7 18.0 38.3 26.9 18.2 24.6 59.4 63.8 32.9 27.3 61.0 57.9 13.1 37.0
MASt3R [duisterhof2024mast3r] 47.5 21.1 23.5 40.2 38.7 38.1 42.2 46.6 66.6 58.0 63.2 40.3 50.4 53.5 15.7 43.1
CUT3R [wang2025continuous3dperceptionmodel] 68.9 50.4 24.7 34.2 37.0 27.6 15.9 21.3 59.1 65.0 65.0 47.5 60.7 66.2 15.1 42.7
MegaSAM [li2024megasam] 73.8 43.9 24.2 29.1 45.3 44.2 11.1 10.2 79.5 82.2 73.8 65.3 71.5 75.8 22.0 50.1
VQAScore
LLaVA-OneVision-7B [li2024llavaonevision] 48.2 13.9 13.0 17.4 24.4 20.8 11.0 14.8 34.5 34.6 17.4 32.3 19.9 21.4 19.4 22.9
LLaVA-Video-7B [zhang2025llavavideo] 56.3 15.7 17.0 19.9 27.9 24.3 16.7 17.4 34.6 37.9 27.7 38.3 16.6 22.4 22.8 26.4
InternVideo2-Chat-8B [wang2024internvideo2] 72.0 19.1 19.9 18.1 18.4 24.1 12.6 10.7 23.3 23.4 17.7 23.5 20.2 16.9 20.8 22.7
Tarsier-Recap-7B [yuan2025tarsier2] 61.5 15.6 26.5 24.4 29.7 22.1 14.8 15.5 23.5 28.1 25.3 22.2 15.7 19.3 31.6 21.6
InternLMXComposer2.5-7B [zhang2024internlmxcomposer25] 50.5 10.9 11.7 10.7 15.0 10.9 12.2 17.0 14.7 14.3 15.1 18.0 12.1 18.6 22.5 17.0
InternVL2.5-8B [chen2025expandingperformance] 69.9 13.3 28.9 26.7 24.1 23.9 19.2 33.1 38.5 31.8 38.7 38.0 11.8 26.1 24.1 30.4
InternVL2.5-26B [chen2025expandingperformance] 65.5 12.2 21.7 24.3 28.0 20.0 22.5 32.5 43.8 39.4 46.2 44.9 14.7 18.7 25.9 30.7
GPT-4o [openai2024gpt4ocard] 68.3 30.1 21.7 39.3 39.1 22.6 43.0 40.5 46.0 43.4 44.9 36.6 24.7 29.6 33.0 37.5
InternVL3-8B [zhu2025internvl3e] 63.0 16.0 19.4 29.9 31.4 28.1 30.4 28.9 42.8 50.8 43.3 37.6 21.9 23.0 20.7 32.4
InternVL3-78B [zhu2025internvl3e] 74.2 18.7 20.2 33.5 34.8 30.3 27.2 34.4 48.6 55.1 49.2 41.5 28.4 25.8 23.3 37.9
Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] 57.7 15.3 19.3 31.4 35.5 28.4 30.7 44.7 64.6 68.7 56.1 35.1 19.4 24.9 20.4 36.9
Qwen2.5-VL-32B [bai2025qwen25vltechnicalreport] 59.3 16.9 20.8 33.1 37.1 30.1 32.3 46.4 66.2 70.2 57.7 36.7 20.8 26.5 21.8 38.4
Qwen2.5-VL-72B [bai2025qwen25vltechnicalreport] 59.7 17.3 21.2 33.5 37.5 30.4 32.7 46.8 66.6 70.7 58.1 37.1 21.2 26.9 22.2 38.8
SOTA Baseline
Cam-Motion-7B [lin2025towards] 86.4 40.5 29.2 50.2 71.3 52.5 57.2 79.6 83.2 87.1 80.1 71.0 33.9 43.1 77.3 63.0
Cam-Motion-32B [lin2025towards] 88.2 42.1 30.8 51.9 73.1 54.1 58.8 81.2 84.7 88.8 81.8 72.7 35.6 44.8 79.2 64.7
Cam-Motion-72B [lin2025towards] 89.4 43.4 32.0 53.1 74.2 55.2 60.0 82.4 86.0 90.1 83.1 73.9 36.8 46.0 80.4 65.9
[…]
As presented in Tab. 4.2, our proposed CamReasoner-7B demonstrates a significant performance leap over a broad spectrum of baseline models, ranging from traditional geometric frameworks to state-of-the-art large vision-language models. Collectively, our model achieves a superior overall accuracy of 78.4%, which not only dramatically outperforms general-purpose benchmarks like GPT-4o and Qwen2.5-VL-72B but also yields a substantial improvement over the task-specific SFT baseline. This performance gap underscores the model’s exceptional mastery of camera-centric reasoning, particularly in complex rotation and translation tasks. Specifically, CamReasoner-7B establishes new SOTA results in challenging categories such as Move Up (84.9%), Tilt Up (87.6%), and Pan Left (86.1%), suggesting that it has developed a robust spatial-temporal awareness capable of decoupling ego-motion from dynamic scene backgrounds. Furthermore, the model exhibits remarkable resilience in traditionally difficult scenarios like Move Out and Zooming, where perspective transformations are most aggressive; for instance, it elevates the performance of the Move Out category to 57.7%, nearly doubling the performance of several competitive baselines. By comparing these results with SFT-only variants, it becomes evident that while supervised learning provides the fundamental paradigm for motion identification, our reinforcement learning phase effectively sharpens the policy’s decision boundaries and maintains high policy entropy during the initial stages to avoid overfitting. Consequently, these results validate that our framework successfully bridges the gap between raw visual perception and high-level camera movement reasoning, providing a robust foundation for automated cinematography analysis.
Table 2: VQA evaluation on the CameraBench. We compare CamReasoner-7B against various state-of-the-art MLLMs across nine fine-grained dimensions with questions on camera movement type in given video clips. Our model consistently achieves the highest accuracy in all categories, demonstrating superior reasoning capabilities in camera movement understanding.
{NiceTabular}
[…]
Steadiness Dynamics Speed Direction Motion Motion Tracking Motion Description Overall
Random Chance 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0
mPLUG-Owl3-7B 53.1 66.4 62.9 49.8 50.4 55.4 54.5 47.1 64.9 57.2
LLaVA-Video-7B [zhang2025llavavideo] 54.8 67.7 58.6 53.4 51.1 56.3 61.3 52.6 69.6 60.2
LLaVA-OneVision-7B [li2024llavaonevision] 55.7 65.3 70.6 54.4 56.8 62.4 62.2 44.4 53.6 58.5
InternVideo2-Chat-8B [wang2024internvideo2] 53.7 65.9 53.0 51.5 51.0 53.5 49.7 52.2 51.8 52.6
Tarsier-Recap-7B [yuan2025tarsier2] 53.1 64.3 51.7 51.0 50.2 52.8 49.0 51.4 51.0 51.8
InternVL2.5-8B [chen2025expandingperformance] 55.8 61.2 58.9 52.6 50.9 59.5 56.5 51.2 51.2 55.8
InternVL2.5-26B [chen2025expandingperformance] 57.6 65.0 62.3 55.1 52.4 61.8 59.8 53.8 54.9 58.6
InternVL3-8B [zhu2025internvl3e] 55.8 61.2 58.9 52.6 50.9 59.5 56.5 51.2 51.2 55.8
InternVL3-78B [zhu2025internvl3e] 57.6 65.0 62.3 55.1 52.4 61.8 59.8 53.8 54.9 58.6
GPT-4o [openai2024gpt4ocard] 57.2 53.9 62.7 59.5 54.6 65.6 53.0 43.1 63.4 60.4
Gemini-2-Flash [comanici2025gemini25] 54.9 47.9 58.0 45.6 42.1 47.6 47.6 40.2 65.3 53.1
Gemini-2.5-Pro [comanici2025gemini25] 59.6 52.5 61.5 50.1 46.8 53.6 50.9 43.9 66.1 56.1
Qwen2.5-VL-3B [bai2025qwen25vltechnicalreport] 59.3 58.3 60.6 54.0 51.7 55.7 57.6 50.0 66.7 59.2
Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] 56.8 60.6 62.9 56.9 54.0 58.4 64.2 53.6 64.4 60.3
Qwen2.5-VL-72B [bai2025qwen25vltechnicalreport] 57.7 62.1 71.0 57.8 53.2 62.4 59.2 50.4 68.6 60.3
[…]
These findings are further corroborated by the VQA evaluation results presented in Tab. 4.2. CamReasoner-7B (Ours) maintains its dominance across a diverse array of evaluative dimensions, achieving a superior overall average accuracy of 74.5%. Our model establishes new performance ceilings in critical categories such as Motion & Steadiness (78.0%), Scene Dynamics (77.0%), and Shot Tracking (71.3%). Notably, the model demonstrates exceptional proficiency in identifying complex scenarios, reaching 82.6% in Complex Description and 60.7% in Confusable Motion, significantly outperforming previous state-of-the-art results. Furthermore, it shows high precision in logical reasoning tasks, as evidenced by its performance in Has Motion (71.8%) and Only Motion (60.6%) categories. This consistent outperformance across both binary classification and multidimensional VQA tasks reinforces the model’s robust capacity for high-fidelity camera movement understanding.
Figure 3: Qualitative results across four typical camera movements. For each case, we visualize the temporal frame sequence alongside the CamReasoner-7B response. The model demonstrates robust spatial reasoning by generating detailed <observation> of visual cues and a logical <think> process to accurately identify the movement and provide the final <answer>.
4.3 Qualitative Results
The qualitative results in Fig. 3 substantiate the superior performance of CamReasoner-7B in camera movement reasoning. Through a structured <observation>-<think>-<answer>pipeline, the model effectively translates fine-grained visual cues into logical deductions.
During the observation phase, the model performs a detailed temporal scan to capture subtle shifts in subject-background relations. This granular perception allows it to decouple ego-motion from environmental noise, explicitly describing transitions between visual landmarks. In the subsequent thinking stage, CamReasoner engages in deductive verification grounded in cinematographic principles. By cross-referencing perceived displacements with the absence of confounding rotations, the model demonstrates that it has internalized robust spatial-temporal paradigms rather than relying on superficial statistical correlations.
Finally, the alignment between these textual reasoning chains and the high accuracy underscores the efficacy of our training strategy. By refining a foundational reasoning checkpoint through reinforcement learning, we successfully bridge the gap between low-level visual perception and high-level semantic understanding.
4.4 Training Curves
Figure 4: Training curves for SFT and RL phases. The top row shows the loss and grad_norm during the SFT process; the bottom row visualizes the convergence of accuracy, format, and overall rewards during RL training.
The training dynamics of the SFT phase exhibit robust convergence and stability, as evidenced by the synchronized progression of loss, gradient norm, and learning rate in Figure 4. Specifically, the training loss undergoes a rapid initial decline and follows a distinct step-wise reduction at each epoch boundary, eventually stabilizing at approximately 0.05, which indicates that the model has effectively captured the underlying distribution of camera movement reasoning tasks. Throughout this process, the gradient norm remains within a well-behaved range between 2.0 and 4.0, ensuring a stable optimization landscape under the peak learning rate of $1.0\times 10^{-5}$ and its subsequent cosine decay. Notably, we select the checkpoint from the conclusion of the first epoch as the initialization for the subsequent RL stage; at this juncture, the model has successfully acquired the fundamental paradigms of spatial reasoning and the <observation>-<think>-<answer>output format, yet retains sufficient policy entropy by avoiding the potential overfitting characteristic of later training stages.
The RL phase, initialized from the first-epoch SFT checkpoint, further refines the model’s policy to achieve rigorous logical alignment. As illustrated in Figure 4, the training dynamics demonstrate a consistent upward trend across all reward metrics. Specifically, the reward/format metric experiences a rapid surge within the first 20 steps, quickly stabilizing near a perfect score of 1.0. This indicates that the model swiftly masters the <observation>-<think>-<answer>structural constraint, even under the exploration pressure of the RL environment. Concurrent with format stabilization, the reward/accuracy exhibits a steady and robust climb from approximately 0.68 to over 0.85, proving that the GRPO-based optimization effectively sharpens the model’s camera-centric reasoning boundaries. The synchronized growth of the reward/overall metric, which eventually plateaus at a high equilibrium, confirms the successful balance between linguistic consistency and spatial-temporal accuracy. As shown in the qualitative results, this optimized policy enables CamReasoner-7B to execute meticulous deductive chains—such as explicitly verifying the absence of confounding motions in complex Truck or Pedestal shots—thereby bridging the gap between raw perception and logical cinematic analysis.
4.5 Ablation Study
Table 3: Ablation of training stages. SFT enables structured reasoning, while RL further polishes performance.
Model SFT RL Binary Avg. VQA Avg.
Base (Qwen2.5-VL-7B) 35.8 60.3
CamReasoner-SFT 68.0 69.5
CamReasoner-7B 78.4 74.5
Table 4: Ablation study of the reward weight $\lambda$. A balanced weight of $\lambda=0.1$ achieves the optimal performance.
$\lambda$ Value 0.0 0.1 0.3 0.5
Binary Avg. 76.2 78.4 75.9 71.4
VQA Avg. 71.8 74.5 72.1 68.3
We evaluate the impact of training stages and the reward weight $\lambda$ on CameraBench. As shown in Table 3, the base model shows limited capability in interpreting complex cinematographic motions. Our SFT stage provides a foundational boost by aligning the model with the O-T-A reasoning paradigm, enabling it to decouple camera ego-motion from scene dynamics. The subsequent RL phase yields an additional 10.4% improvement in binary accuracy, demonstrating its effectiveness in optimizing decision-making and enforcing logical consistency across long-range temporal sequences.
Furthermore, Table 4 analyzes the sensitivity of $\lambda$, which balances accuracy ($r_{acc}$) and format ($r_{fmt}$) rewards. Performance peaks at $\lambda=0.1$, suggesting that while a small format reward is crucial for stabilizing the reasoning structure and ensuring parseable outputs, over-prioritizing formatting at higher values can detract from deep reasoning by constraining the model’s expressive capacity. This indicates that once the structural paradigm is established, the reinforcement signal should primarily focus on the correctness of the spatial-temporal deductions.
5 Conclusions
In this paper, we introduced CamReasoner, a framework that reformulates camera movement understanding as a structured ego-motion inference task. By employing the O-T-A paradigm, the model explicitly decouples camera motion from complex backgrounds through spatial-temporal reasoning. Our construction of a Large-scale Inference Trajectory Suite, comprising 18k SFT and 38k RL samples, provides a foundational resource for teaching models cinematic logic. Experimental results show that CamReasoner-7B establishes a new state-of-the-art, achieving 78.4% accuracy in binary classification and outperforming GPT-4o by over 40%. Furthermore, our pioneering use of Reinforcement Learning for logical alignment effectively mitigates hallucinations in challenging scenarios like Confusable Motion. By bridging visual perception and structured deduction, CamReasoner offers a powerful foundation for future advancements in automated film analysis and controllable video synthesis.